OpenStudy (anonymous):

establish an equation to determine revenue and total cost from a fixed cost of $55, sell price of $15 and labor cost $4. Interpret your result from each.

OpenStudy (callisto):

total cost = fixed cost + labour cost = 55+4 =59 (just sum up the 'cost') revenue = selling price - total cost (sub the values) i wonder why it turns into a loss rather than revenue :S

OpenStudy (anonymous):

thats what I can't figure out. so the revenue will be the 59

OpenStudy (anonymous):

cuz i have to decide how many pizzas have to be sold to break even

OpenStudy (callisto):

i think that is something about accounting?

OpenStudy (anonymous):

what happened to the$15

OpenStudy (callisto):

wait!

OpenStudy (anonymous):

i have no idea on what i am doing so i am relying on your help

OpenStudy (callisto):

is labour cost a fixed cost or variable cost?

OpenStudy (anonymous):

i have to sell pizzas for 15. fixed cost is 55 and labor cost is 4

OpenStudy (anonymous):

i dont have anything on unit contribution margin. i just know how many pizzas must be sold to break even. Where do I include the price of the pizza for $15....you don't mention that again

OpenStudy (callisto):

oh, at break-even point, sale revenue=Fixed cost + variable cost in your case, let x be the no. of pizza sold that would be sale revenue=Fixed cost + labour cost 15X = 55+ 4X X=5 done!

OpenStudy (callisto):

ignore the words 'that would be'

OpenStudy (anonymous):

ok now how would i do that if i have to determine profit of selling 500 pizzas. interpret result

OpenStudy (callisto):

but i don't really know if labour cost depends on the no. of unit you produced

OpenStudy (anonymous):

the revenue would be the 59 and the total cost is 55

OpenStudy (anonymous):

idk either, i will have to talk to teacher again...i just wanted some help getting this done because it is due tmrw and i don't know what i am doing

OpenStudy (callisto):

the revenue would be the 59, is it given in the question or the answer?

OpenStudy (anonymous):

it is in the answer

OpenStudy (anonymous):

so just to make sure my break even is a profit of 5 pizzas or negative 5

OpenStudy (callisto):

break even actually means that you do not have loss/gain when you produce that amount of product

OpenStudy (callisto):

i just wonder why the revenue is 59, it's true only when you produce one pizza / when the labour cost is the fixed cost

OpenStudy (callisto):

perhaps you can do both cases. above case is when labour cost is a variable cost

OpenStudy (anonymous):

establish an equation to determine total cost? So what would the equation be for selling 500 pizzas

OpenStudy (callisto):

for labour cost being the fixed cost, let x be the no. of pizza sold that would be sale revenue=Fixed cost + labour cost 15X = 55+ 4 X=3.933

OpenStudy (callisto):

total cost of selling 500 pizzas = 55+4x500 = 2055 (only when labour cost is variable cost, if it is the fixed one, the total cost won't change!)

OpenStudy (anonymous):

can you show me how you came up with that x = 3.933 step by step by any chance

OpenStudy (anonymous):

ok i apologize i am asking you so many questions, how many pizzas need to be sold to make a profit of $1100.

OpenStudy (callisto):

15X = 55+ 4 15X = 59 X= 59/15 X=3.933

OpenStudy (anonymous):

thank you for being patient with me

OpenStudy (callisto):

can you try first, it's similar but using the formula profit = sale revenue - variable cost - fixed cost

OpenStudy (anonymous):

2055= 59 - 4-55

OpenStudy (anonymous):

Actually, Cost = 55 + 4q

OpenStudy (anonymous):

Several solvers already responded to your previous posts! It already solved!

OpenStudy (anonymous):

i am asking different questions

OpenStudy (anonymous):

i have only got one reply back that i can see

OpenStudy (anonymous):

It's exactly the same concept!

OpenStudy (anonymous):

@Callisto where it confuses you is the labor cost, it should be 4x!

OpenStudy (anonymous):

i may be asking the same question because i don't get the concept.

OpenStudy (callisto):

thanks chlorophyll, i thought so :)

OpenStudy (anonymous):

Revenue = unit cost * quantity

OpenStudy (anonymous):

Do you follow at this point?

OpenStudy (callisto):

the concept is that: sales revenue = selling price x no. of unit sold total cost = total fixed cost + total variable cost , total variable cost = variable cost per unit x no. of unit sold to calculate profit, you should use sales revenue - total cost ie. profit = revenue - cost at break even point, net profit =0 therefore sales revenue = total cost

OpenStudy (anonymous):

@babybluz24149 we'll go step by step!

OpenStudy (anonymous):

no im sorry, i hate algebra

OpenStudy (anonymous):

Don't think algebra, just like you go sale ice cream!

OpenStudy (anonymous):

ok i like money i just don't understand the x and the y concept

OpenStudy (anonymous):

for example: 1 ice cream costs $2

OpenStudy (anonymous):

then 5 ice creams will be ...

OpenStudy (anonymous):

10

OpenStudy (anonymous):

See, when you multiply quantity *unit price you get REVENUE

OpenStudy (anonymous):

Is it okie here?

OpenStudy (anonymous):

ok i can understand that way. i guess just the way its worded and i have to put in an equation that i have no idea on what i am doing

OpenStudy (anonymous):

Trust me, it's not hard at all :)

OpenStudy (anonymous):

then assume q is 10 ice cream!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then you set your price $15 per ice cream!

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

Then, tell me what is your REVENUE here?

OpenStudy (anonymous):

150

OpenStudy (anonymous):

so that is how i get stuck writing out the equation with an x

OpenStudy (anonymous):

Now we different numbers, any number, not just the quantity 10

OpenStudy (anonymous):

What is the revenue with any quantity?

OpenStudy (anonymous):

@Callisto did you have solution with your post yet?

OpenStudy (callisto):

huh?

OpenStudy (anonymous):

@babybluz24149 what is your new revenue with any quantity?

OpenStudy (anonymous):

i didn't the one for how many pizzas need to be sold to make a profit of 1,100

OpenStudy (callisto):

don't think about the pizza first! think about ice cream!

OpenStudy (anonymous):

You need to get your mind fixed with the revenue here, respond to me!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what is REVENUE with any number of ice cream?

OpenStudy (anonymous):

1100* 10=11000

OpenStudy (callisto):

think about the ice cream first :S

OpenStudy (anonymous):

Thanks, Callisto :)

OpenStudy (anonymous):

Any number means no number ==> X

OpenStudy (anonymous):

@babybluz24149 WHERE ARE YOU?

OpenStudy (anonymous):

im still here

OpenStudy (anonymous):

sorry

OpenStudy (callisto):

remember you sell 1 ice cream for $15 when you sell 1, you get $15 when you sell 2, you get 2 x $15, that is $30 when you sell 10, you said you'd get $150

OpenStudy (anonymous):

right

OpenStudy (callisto):

so how did you work out $150?

OpenStudy (anonymous):

Now we're tired of repeated, so we don't want number any more!

OpenStudy (anonymous):

multiplied the number of ice cream by the price

OpenStudy (anonymous):

Yep, that called REVENUE

OpenStudy (anonymous):

ok i got it

OpenStudy (callisto):

hurray!!!

OpenStudy (callisto):

what did you get then?

OpenStudy (anonymous):

lol i know how to multiply its what to do and when

OpenStudy (anonymous):

Here you already know unit price is $15/ ice cream!

OpenStudy (anonymous):

right

OpenStudy (anonymous):

the amount is called the revenue

OpenStudy (anonymous):

Unfortunately, you want to find the number, quantity of ice cream

OpenStudy (anonymous):

ok

OpenStudy (callisto):

chlorophyll, you haven't finished the cost part!

OpenStudy (anonymous):

=> REVENUE = 15 * .....

OpenStudy (anonymous):

@callisto, Have I type the first sentence is "step by step"

OpenStudy (callisto):

continue then

OpenStudy (anonymous):

@babybluz24149 , you still haven't answered what your REVENUE with $15 per ice cream!